On average a certain computer part lasts ten years. The length of time the computer part lasts is exponentially distributed .

i)What is the probability that a computer part lasts more than 7 years?

ii) What is the probability that a computer part lasts between 9 and 11 years?

iii)Eighty percent of computer parts lasts for how long?


1
Expert's answer
2022-01-04T08:34:05-0500

for exponential distribution:

"P(X<x)=1-e^{-mx}"

we have: "m=1\/10"


a)

"P(X>7)=e^{-0.1\cdot7}=0.4966"


b)

"P(9<X<11)=P(X<11)-P(X<9)=e^{-0.9}-e^{-1.1}=0.0737"


c)

"1-e^{-0.1x}=0.8"

"x=-ln0.2\/0.1=16.1" years